Square Manufacturing is considering investing in a robotics manufacturing line. Installation of the line will cost an estimated $10.2 million. This amount must be paid immediately even though construction will take three years to complete (years 0, 1, and 2). Year 3 will be spent testing the production line and, hence, it will not yield any positive cash flows. If the operation is very successful, the company can expect after-tax cash savings of $7.2 million per year in each of years 4 through 7. After reviewing the use of these systems with the management of other companies, Square’s controller has concluded that the operation will most probably result in annual savings of $5.4 million per year for each of years 4 through 7. However, it is entirely possible that the savings could be as low as $3.0 million per year for each of years 4 through 7. The company uses a 16 percent discount rate. Use Exhibit A.8.

Required:

Compute the NPV under the three scenarios. (Round PV factor to 3 decimal places. Enter your answers in thousands of dollars. Negative amounts should be indicated by a minus sign.)
Answer #1

Here's the calculation for all 3 scenarios:

Discount Rate (i) 16%
Scenario 1:
Year (n) 1 2 3 4 5 6 7
After tax cash savings (FV) 0 0 0 7200000 7200000 7200000 7200000
PV Factor =" (1+i)" to the power "n" (rounded off to 3 decimal places) 1.16 1.346 1.561 1.811 2.1 2.436 2.826
NPV Formulae = FV/PV factor (Value in $'000s) 0 0 0 3975.704 3428.5714 2955.665 2547.771
Thus, NPV = sum of all 7 years (Value in $'000s) $12,907.71
Scenario 2:
Year (n) 1 2 3 4 5 6 7
After tax cash savings (FV) 0 0 0 5400000 5400000 5400000 5400000
PV Factor =" (1+i)" to the power "n" (rounded off to 3 decimal places) 1.16 1.346 1.561 1.811 2.1 2.436 2.826
NPV Formulae = FV/PV factor (Value in $'000s) 0 0 0 2981.778 2571.4286 2216.749 1910.828
Thus, NPV = sum of all 7 years (Value in $'000s) $9,680.78
Scenario 3:
Year (n) 1 2 3 4 5 6 7
After tax cash savings (FV) 0 0 0 3000000 3000000 3000000 3000000
PV Factor =" (1+i)" to the power "n" (rounded off to 3 decimal places) 1.16 1.346 1.561 1.811 2.1 2.436 2.826
NPV Formulae = FV/PV factor (Value in $'000s) 0 0 0 1656.543 1428.5714 1231.527 1061.571
Thus, NPV = sum of all 7 years (Value in $'000s) $5,378.21

> Why would you divide the FV by the PV instead of multiplying the present value factor from the Present Value of $1 table using 16% and 7 periods

Christian Ming Xiong Tue, Oct 19, 2021 7:41 PM